OpenStudy (anonymous):

Let f(x) = x + 6 and g(x) = x – 5. What is f(x) • g(x)? Answer: A) x2 – 6x + 1 B) x2 – x + 30 C) x2 + 6x + 30 D) x2 + x – 30 help me please?:)

jimthompson5910 (jim_thompson5910):

f(x)*g(x) (x+6)(x-5) x^2-5x+6x-30 x^2+x-30 So the answer is choice D

OpenStudy (anonymous):

Yay! thank you! What about this one? Let f(x) = x – 1 and g(x) = 2x + 3. What is f(g(x)) ? Answer: A) x + 4 B) 3x + 2 C) 2x + 2 D) 2x + 1 :)))

jimthompson5910 (jim_thompson5910):

f(x) = x-1 f(g(x)) = g(x)-1 f(g(x)) = (2x+3) - 1 f(g(x)) = 2x + 2 so choice C

OpenStudy (anonymous):

thank you<33

jimthompson5910 (jim_thompson5910):

np, glad to be of help